OpenStudy (anonymous):

Temperature and volume formula V(T)= 0.147T+40 calculates the volume V in cubic inches of a sample of gas at temperature T in degrees celsius a. evaluate V(0) and interpet the result b. if the temperature increases by 10 degree celsius by how much does the volume increase? c. what is the volume of gas when the temperature is 100 degrees celsius?

OpenStudy (anonymous):

Please Help?????????

OpenStudy (anonymous):

What does V equal when T=0?

OpenStudy (anonymous):

if T=0 then imput it in the equation 0.147(0)+40=V

OpenStudy (anonymous):

so V would equal 40

OpenStudy (anonymous):

a has been done by @mads4566 for b let v1=0.147T+40 v2=0.147(T+10)+40 now subtract v1 from v2 and find out the change in volume

OpenStudy (anonymous):

for c simply plug in t=100 in V(t)=0.147T+40 and u get the answer

OpenStudy (anonymous):

Its easier to use v0 and v1 where v0 is when T=0 and v1 is when T=10

OpenStudy (anonymous):

*T=100 instead of t

OpenStudy (anonymous):

ok I think I have it, ty so much everyone....wish i had you by my side, I need ALOT of help this weekend to play catch up from being in the hospital, finals are monday and I need all of this turned in :(

OpenStudy (anonymous):

I wish i could give all of you medals :(